OpenStudy (anonymous):

Factor x3 + 2x2 + 5x + 10

OpenStudy (anonymous):

x is common x(x^2+2x+5+10)=0 so x=0 x^2+2x+15 can you continue?

OpenStudy (anonymous):

x isn't actually common to the 10 so you cannot factor it out like that.

myininaya (myininaya):

x^2(x+2)+5(x+2)=(x+2)(x^2+5)